OpenStudy (james1107):

Find the midpoint of the line segment whose endpoints are (2.6, 5.1) and (3, 4.7). (0.2, 0.2) (1.45, 4.9) (2.8, 4.9)

OpenStudy (james1107):

@Keigh2015

OpenStudy (james1107):

@janu16

OpenStudy (janu16):

got it

OpenStudy (james1107):

no not at all

OpenStudy (janu16):

no i said i got it its up there

OpenStudy (janu16):

The midpoint of the line segment whose endpoints are (2.6, 5.1) and (3, 4.7) is (2.8, 4.9). M = [(x1+x2)/2, (y1+y2)/2]; M = [(2.6+3)/2, (5.1+4.7)/2]; M = (5.6/2, 9.8/2); M = (2.8, 4.9)

OpenStudy (janu16):

right there

OpenStudy (janu16):

its all jumbled up but do you see it?

OpenStudy (james1107):

you are confusing me

OpenStudy (keigh2015):